OpenStudy (anonymous):

List the discontinuities for the function f(x) = cot( 2x/3) a: n(3pi/2) when n is an integer b: n(pi/2) when n is an integer c: n(pi/3) when n is an integer d: there are no discontinuities

OpenStudy (anonymous):

\[f(x) = \cot(2x/3) = \frac{ \cos(\frac{ 2x }{ 3 }) }{ \sin (\frac{ 2x }{ 3 }) }\] the discontinuity happens when sin(2x/3) = 0 if you make x = 3pi/2 you find that sin(2*(3pi/2)/3) = sin(pi) = 0

OpenStudy (anonymous):

I found (a) as answer

OpenStudy (anonymous):

thanks!